1 g of NAOH is added to 30 cm3 of 1 M HCL. How many cubic cm of 0.1 KOH will be needed to neutralize excess acid
1
Expert's answer
2020-07-14T15:19:01-0400

Molecular weight of NaOH is 40 g/mol

Moles of NaOH added = 1/40 mol = 0.025 mol

Strength of NaOH is = (0.025 ×1000)/30 = 0.8333 M

Excess acid presence = (1-0.833) M = 0.1667 M in 30 cm3

V1S1=V2S2

V1 = (0.1667×30 )/0.1 = 50.01 cm3

50.01 cm3 KOH will be needed to neutralize excess acid.